# A Survey of NL2SQL with Large Language Models: Where are we, and where are we going?

### v4

@@ Abstract
Translating users’ natural language queries (NL) into SQL queries (i.e., NL2SQL, a.k.a., Text-to-SQL) can significantly reduce barriers to accessing relational databases and support various commercial applications. The performance of NL2SQL has been greatly enhanced with the emergence of Large Language Models (LLMs). In this survey, we provide a comprehensive review of NL2SQL techniques powered by LLMs, covering its entire lifecycle from the following four aspects: (1) Model: NL2SQL translation techniques that tackle not only NL ambiguity and under-specification, but also properly map NL with database schema and instances; (2) Data: From the collection of training data, data synthesis due to training data scarcity, to NL2SQL benchmarks; (3) Evaluation: Evaluating NL2SQL methods from multiple angles using different metrics and granularities; and (4) Error Analysis: analyzing NL2SQL errors to find the root cause and guiding NL2SQL models to evolve. Moreover, we offer a rule of thumb for developing NL2SQL solutions. Finally, we discuss the research challenges and open problems of NL2SQL in the LLMs era.
@@ I Introduction
Natural Language to SQL (i.e., nl2sql, a.k.a., Text-to-SQL), which converts a natural language query (nl) into an sql query, is a key technique toward lowering the barrier to accessing relational databases  [1, 2, 3, 4, 5, 6, 7]. This technique supports various important applications such as business intelligence and natural language interfaces for databases, making it a key step toward democratizing data science  [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Recent advancements in language models have significantly extended the frontiers of research and application in nl2sql. Concurrently, the trend among database vendors to offer nl2sql solutions has evolved from a mere notion to a necessary strategy [20, 21]. Therefore, we need to understand the fundamentals, techniques, and challenges regarding nl2sql.
@@
In this survey, we will systematically review recent nl2sql techniques through a new framework, as shown in Figure LABEL:fig:overview.
@@
  • NL2SQL with Language Models. We will first review existing nl2sql solutions from the perspective of language models, categorizing them into four major categories (see Figure LABEL:fig:overview(a)). We will then focus on the recent advances in Pre-trained Language Models (PLMs) and Large Language Models (LLMs) for nl2sql.
@@
  • Benchmarks and Training Data Synthesis. Undoubtedly, the performance of PLM- and LLM-based nl2sql models is highly dependent on the amount and quality of the training data. Therefore, we will first summarize the characteristics of existing benchmarks and analyze their statistical information (e.g., database and query complexity) in detail. We will then discuss methods for collecting and synthesizing high-quality training data, emphasizing this as a research opportunity (see Figure LABEL:fig:overview(b)).
@@
  • Evaluation. Comprehensively evaluating nl2sql models is crucial for optimizing and selecting models for different usage scenarios. We will discuss the multi-angle evaluation and scenario-based evaluation for the nl2sql task (see Figure LABEL:fig:overview(c)). For example, we can assess the nl2sql model in specific contexts by filtering benchmarks based on SQL characteristics, nl variants, database domains, and so on.
@@
  • NL2SQL Error Analysis. Error analysis is essential in nl2sql research for identifying limitations and improving the model robustness. We review existing error taxonomies, analyze their limitations, and propose principles for designing comprehensive taxonomies for nl2sql output errors. Using these principles, we create a two-level error taxonomy and utilize it to summarize and analyze nl2sql output errors (see Figure LABEL:fig:overview(d)).
@@
In addition to the above, we will provide practical guidance for developing nl2sql solutions, including a roadmap for optimizing LLMs for nl2sql tasks and a decision flow for selecting nl2sql modules tailored to various nl2sql scenarios. Finally, we will discuss key open problems in the field, such as open nl2sql tasks, cost-effective nl2sql with LLMs, and trustworthy nl2sql solutions.
@@
Differences from Existing Surveys. Our survey distinguishes itself from existing nl2sql surveys [22, 23, 24, 25, 26, 27, 28, 29] and tutorials [30, 31, 32] in five aspects.
@@
  • We systematically review the entire lifecycle of NL2SQL problem, as shown in Figure LABEL:fig:overview. This lifecycle includes various nl2sql translation methodologies powered by language models (Figure LABEL:fig:overview(a)), training data collection and synthesis methods (Figure LABEL:fig:overview(b)), multi-angle and scenarios-based evaluations (Figure LABEL:fig:overview(c)), and nl2sql output error analysis techniques (Figure LABEL:fig:overview(d)).
@@
  • We provide a more detailed and comprehensive summary of the inherent challenges in nl2sql. Additionally, we analyze the technical challenges when developing a robust nl2sql solution for real-world scenarios, which are often overlooked in other surveys.
@@
  • We particularly focus on recent advances in LLM-based nl2sql methods, summarizing key modules and comparing different strategies within this scope. We are the first survey to provide a modular summary of methods and provide detailed analyses for each key module (e.g., database content retrieval).
@@
  • We highlight the importance of evaluating nl2sql methods in a multi-angle way, analyze the key nl2sql error patterns, and provide a two-level error taxonomy.
@@
  • We provide practitioners with a roadmap for optimizing LLMs to nl2sql task and a decision flow for selecting the suitable nl2sql modules for various usage scenarios.
@@
  • NL2SQL with Language Models. We comprehensively review existing nl2sql techniques from a lifecycle perspective (Figure LABEL:fig:overview). We introduce the nl2sql task definition, discuss challenges (Figure 1), provide a taxonomy of nl2sql solutions based on language models (Figure 2), and summarize the key modules of language model-powered nl2sql solutions (Figure 4 and Table I). Next, we elaborate on each module of language model-powered nl2sql methods, including the pre-processing strategies (Section IV), nl2sql translation methods (Section V), and post-processing techniques (Section VI).
  • Benchmarks and Training Data Synthesis. We summarize existing nl2sql benchmarks based on their characteristics (Figure 9). We analyze each benchmark in depth and discuss its pros and cons (Table II). (Section VII)
  • Evaluation and Errors Analysis. We highlight the importance of evaluation in developing practical nl2sql solutions. We review widely used evaluation metrics and toolkits for assessing nl2sql solutions. We provide a taxonomy to summarize typical errors produced by nl2sql methods. (Section VIII)
  • Practical Guidance for Developing NL2SQL Solutions. We provide a roadmap for optimizing existing LLMs to nl2sql tasks (Figure 10(a)). In addition, we design a decision flow to guide the selection of appropriate nl2sql modules for different scenarios (Figure 10(b)).
  • Open Problems in NL2SQL. Finally, we discuss new research opportunities, including the open-world nl2sql problem and cost-effective nl2sql solutions (Section X).
  • NL2SQL Handbook. We maintain an online handbook (https://github.com/HKUSTDial/NL2SQL_Handbook) to help readers stay current with advancements in nl2sql.
II NL2SQL Problem and Background
In this section, we first formalize the definition of the nl2sql task (Section II-A). We then introduce the workflow of how humans perform the nl2sql task (Section II-B) and discuss the key challenges (Section II-C). Finally, we describe the evolution of nl2sql solutions based on the development of language models (Section II-D).
@@ II-A Problem Formulation
Definition 1 (Natural Language to SQL (NL2SQL)).
Natural Language to SQL (nl2sql), also known as Text-to-SQL, is the task of converting natural language queries (nl) into corresponding SQL queries (sql) that can be executed on a relational database (db). Specifically, given an nl and a db, the goal of nl2sql is to generate an sql that accurately reflects the user’s intent and returns the appropriate results when executed on the database.
@@
Figure 1: Examples of the nl2sql Task and Its Challenges.
II-B NL2SQL Human Workflow
When professional users (e.g., DBAs) perform the nl2sql task, they first interpret the nl question, examine the database schema and contents, and then construct the corresponding sql based on their SQL expertise. Below, we outline this process in detail, as illustrated in Figure 1(a).
@@ II-B NL2SQL Human Workflow
Note that Step-2 aligns with the concepts of schema linking, database content retrieval, and additional information acquisition in recent nl2sql solutions powered by language models (please refer to Figure 4 for more details).
Step-3: Writing SQL based on NL and DB Understanding: Finally, the DBA writes the corresponding sql based on the insights gained in Steps-1 and -2. This process, known as “nl2sql Translation”, relies heavily on the DBA’s SQL expertise. However, this process can be very challenging due to the ambiguity of the nl or the complexity of the database. For example, as shown in Figure 1(a), despite understanding the need to link the Customer and Book tables, one must be familiar with the usage and norms of employing either a natural join or a subquery. In addition, there may be multiple possible sql queries because “genres” can refer to either “LiteraryGenre” or “SubjectGenre”.
Takeways. From the above steps, we intuitively identify three inherent challenges in the nl2sql task: the uncertainty of the natural language, the complexity of the database, and the translation from the “free-form” natural language queries to the “constrained and formal” sql queries.
II-C NL2SQL Task Challenges
In this section, we will first discuss the fundamental challenges of the nl2sql task. We will then analyze the technical challenges, i.e., the challenges we face when developing a strong nl2sql solution in real-world scenarios.
C1: Uncertain Natural Language Query. Natural language queries often contain uncertainties due to ambiguity and underspecification [33]. In nl2sql tasks, the challenges related to nl can be summarized as follows:
  • Lexical Ambiguity: This occurs when a single word has multiple meanings. For example, the word “bat” can refer to an animal or a baseball bat (noun) or the action of swinging (verb).
@@ II-C NL2SQL Task Challenges
Figure 2: The Evolution of NL2SQL Solutions from the Perspective of Language Models.
C2: Complex Database and Dirty Content. The nl2sql task requires a deep understanding of the database schema, including table names, columns, relationships, and data attributes. The complexity of modern schemas and large data volumes make this task especially challenging.
  • Complex Relationships Among Tables: Databases often contain hundreds of tables with complex interrelationships. nl2sql solutions must accurately comprehend and leverage these relationships when generating sql queries.
  • Ambiguity in Attributes and Values: Ambiguous values and attributes in a database can complicate nl2sql systems’ ability to identify the correct context.
  • Domain-Specific Schema Designs: Different domains often have unique database designs and schema patterns. The variations in schema design across domains make it difficult to develop a one-size-fits-all nl2sql solution.
@@
C3: NL2SQL Translation. The nl2sql task differs from the compilation of a high-level programming language to a low-level machine language, as it usually has a one-to-many mapping between the input nl and output sql queries. Specifically, the nl2sql task faces several unique challenges:
@@
  • Database Schema Dependency: The nl2sql translation is highly dependent on the underlying database schema. As shown in Figure 1 (a) and (b), the same nl may produce different sql queries based on schema variations. This requires nl2sql models to bridge gaps between training data and real-world schema differences.
Beyond the intrinsic challenges of the nl2sql task, developers must also overcome several technical obstacles to build reliable and efficient nl2sql systems, as discussed below.
C4: Technical Challenges in Developing NL2SQL Solutions. Developing robust nl2sql solutions requires addressing several key technical challenges, including:
  • Cost-effective Solution: Deploying nl2sql models, particularly those using large language models, demands significant resources, such as hardware and/or API costs. Achieving an optimal balance between model performance and cost efficiency remains a crucial challenge.
@@
  • SQL Efficiency: The sql generated by nl2sql models must be both correct and optimized for performance. This includes optimizing join operations, index usage, and query structures. Efficient queries reduce database load, improving system responsiveness and throughput.
  • Insufficient and Noisy Training Data: High-quality nl2sql training data is challenging to obtain. Public datasets are often limited and may include noisy annotations, affecting model performance [34]. Annotation requires database expertise, increasing costs, and the complexity of nl2sql tasks often leads to errors.
  • Trustworthiness and Reliability: nl2sql models must be trustworthy and reliable, consistently producing accurate results across diverse datasets and scenarios. Trustworthiness requires transparency, allowing users to understand and verify the generated sql.
@@ II-D Challenges Solving with Large Language Models
Difficulty Levels. We categorize the difficulty of nl2sql into five levels, each addressing specific hurdles, as shown in Figure 2(a). The first three levels cover challenges that have been or are currently being addressed, highlighting the gradual progress in nl2sql capabilities. The fourth level includes challenges that are the focus of current LLM-based solutions, while the fifth level represents future challenges, showing our vision for nl2sql advancements over the next five years.
The Evolution of nl2sql Solutions. The development of nl2sql solutions, illustrated in Figure 2(b), progresses through four distinct stages: the rule-based stage, the neural network-based stage, the PLM-based stage, and the LLM-based stage. At each stage, we analyze shifts in target users, i.e., from experts to broader user groups, and the extent to which various nl2sql challenges are addressed.
Figure 3: The Categorization of PLM and LLM in NL2SQL.

### v5

# A Survey of Text-to-SQL in the Era of LLMs: Where are we, and where are we going?

@@ Abstract
Translating users’ natural language queries (NL) into SQL queries (i.e., Text-to-SQL, a.k.a. NL2SQL) can significantly reduce barriers to accessing relational databases and support various commercial applications. The performance of Text-to-SQL has been greatly enhanced with the emergence of Large Language Models (LLMs). In this survey, we provide a comprehensive review of Text-to-SQL techniques powered by LLMs, covering its entire lifecycle from the following four aspects: (1) Model: Text-to-SQL translation techniques that tackle not only NL ambiguity and under-specification, but also properly map NL with database schema and instances; (2) Data: From the collection of training data, data synthesis due to training data scarcity, to Text-to-SQL benchmarks; (3) Evaluation: Evaluating Text-to-SQL methods from multiple angles using different metrics and granularities; and (4) Error Analysis: analyzing Text-to-SQL errors to find the root cause and guiding Text-to-SQL models to evolve. Moreover, we offer a rule of thumb for developing Text-to-SQL solutions. Finally, we discuss the research challenges and open problems of Text-to-SQL in the LLMs era.
@@ I Introduction
Natural Language to SQL (i.e., Text-to-SQL), which converts a natural language query (nl) into an sql query, is a key technique toward lowering the barrier to accessing relational databases  [1, 2, 3, 4, 5, 6, 7]. This technique supports various applications such as business intelligence and natural language interfaces for databases, making it a key step toward democratizing data science  [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Recent advancements in language models have significantly extended the frontiers of research and application in Text-to-SQL. Concurrently, the trend among database vendors to offer Text-to-SQL solutions has evolved from a mere notion to a necessary strategy [19, 20]. Therefore, we need to understand the fundamentals, techniques, and challenges regarding Text-to-SQL.
@@
In this survey, we systematically review recent Text-to-SQL techniques through a new framework, as shown in Figure LABEL:fig:overview.
@@
  • Text-to-SQL with Language Models. We will first review existing Text-to-SQL solutions from the perspective of language models, categorizing them into four major categories (see Figure LABEL:fig:overview(a)). We will then focus on the recent advances in Pre-trained Language Models (PLMs) and Large Language Models (LLMs) for Text-to-SQL.
@@
  • Benchmarks and Training Data Synthesis. Undoubtedly, the performance of PLM- and LLM-based Text-to-SQL models is highly dependent on the amount and quality of the training data. Therefore, we will first summarize the characteristics of existing benchmarks and analyze their statistical information (e.g., database complexity) in detail. We will then discuss methods for collecting and synthesizing high-quality training data, emphasizing this as a research opportunity (see Figure LABEL:fig:overview(b)).
@@
  • Evaluation. Comprehensively evaluating Text-to-SQL models is crucial for optimizing and selecting models for different usage scenarios. We will discuss the multi-angle evaluation and scenario-based evaluation for the Text-to-SQL task (see Figure LABEL:fig:overview(c)). For example, we can assess the Text-to-SQL model in specific contexts by filtering benchmarks based on SQL characteristics, nl variants, database domains, and so on.
@@
  • Text-to-SQL Error Analysis. Error analysis is essential in Text-to-SQL research for identifying limitations and improving the model robustness. We review existing error taxonomies, analyze their limitations, and propose principles for designing comprehensive taxonomies for Text-to-SQL output errors. Using these principles, we create a two-level error taxonomy and utilize it to summarize and analyze Text-to-SQL output errors (see Figure LABEL:fig:overview(d)).
@@
In addition to the above, we will provide practical guidance for developing Text-to-SQL solutions, including a roadmap for optimizing LLMs for Text-to-SQL tasks and a decision flow for selecting Text-to-SQL modules tailored to various Text-to-SQL scenarios. Finally, we will discuss key open problems in the field, such as open Text-to-SQL tasks, cost-effective Text-to-SQL with LLMs, and trustworthy Text-to-SQL solutions.
@@
Differences from Existing Surveys. Our survey distinguishes itself from existing Text-to-SQL surveys [21, 22, 23, 24, 25, 26, 27, 28] and tutorials [29, 30, 31] in five aspects.
@@
  • We systematically review the entire lifecycle of Text-to-SQL problem, as shown in Figure LABEL:fig:overview. This lifecycle includes various Text-to-SQL translation methodologies powered by language models (Figure LABEL:fig:overview(a)), training data collection and synthesis methods (Figure LABEL:fig:overview(b)), multi-angle and scenarios-based evaluations (Figure LABEL:fig:overview(c)), and Text-to-SQL error analysis techniques (Figure LABEL:fig:overview(d)).
@@
  • We provide a more detailed and comprehensive summary of the inherent challenges in Text-to-SQL. Additionally, we analyze the technical challenges when developing a robust Text-to-SQL solution for real-world scenarios, which are often overlooked in other surveys.
@@
  • We particularly focus on recent advances in LLM-based Text-to-SQL methods, summarizing key modules and comparing different strategies within this scope. We are the first survey to provide a modular summary of methods and provide detailed analyses for each key module (e.g., database content retrieval).
@@
  • We highlight the importance of evaluating Text-to-SQL methods in a multi-angle way, analyze the Text-to-SQL error patterns, and provide a two-level error taxonomy.
@@
  • We provide practitioners with a roadmap for optimizing LLMs to Text-to-SQL and a decision flow for selecting the suitable Text-to-SQL modules for various scenarios.
@@
  • Text-to-SQL with Language Models. We comprehensively review existing Text-to-SQL techniques from a lifecycle perspective (Figure LABEL:fig:overview). We introduce the Text-to-SQL task definition, discuss challenges (Figure 1), provide a taxonomy of Text-to-SQL solutions based on language models (Figure 2), and summarize the key modules of language model-powered Text-to-SQL solutions (Figure 4 and Table I). Next, we elaborate on each module of language model-powered Text-to-SQL methods, including the pre-processing strategies (Section IV), Text-to-SQL translation methods (Section V), and post-processing techniques (Section VI).
  • Benchmarks and Training Data Synthesis. We summarize existing Text-to-SQL benchmarks based on their characteristics (Figure 9). We analyze each benchmark in depth and discuss its pros and cons (Table II). (Section VII)
  • Evaluation and Errors Analysis. We highlight the importance of evaluation in developing practical Text-to-SQL solutions. We review widely used evaluation metrics and toolkits for assessing Text-to-SQL solutions. We provide a taxonomy to summarize typical errors produced by Text-to-SQL methods. (Section VIII)
  • Practical Guidance for Developing Text-to-SQL Solutions. We provide a roadmap for optimizing existing LLMs to Text-to-SQL tasks (Figure 10(a)). In addition, we design a decision flow to guide the selection of appropriate modules for different scenarios (Figure 10(b)).
  • Open Problems in Text-to-SQL. We analyze the limitations of LLM-based methods and discuss new research opportunities, including the open-world Text-to-SQL problem and cost-effective solutions (Section X).
  • Text-to-SQL Handbook. We maintain an online handbook (https://github.com/HKUSTDial/NL2SQL_Handbook) to help readers stay current with Text-to-SQL advancements.
II Text-to-SQL Problem and Background
In this section, we first formalize the definition of the Text-to-SQL task (Section II-A). We then introduce the workflow of how humans perform the Text-to-SQL task (Section II-B) and discuss the key challenges (Section II-C). Finally, we describe the evolution of Text-to-SQL solutions based on the development of language models (Section II-D).
@@ II-A Problem Formulation
Definition 1 (Natural Language to SQL (Text-to-SQL)).
Natural Language to SQL (Text-to-SQL), also known as NL2SQL, is the task of converting natural language queries (nl) into corresponding SQL queries (sql) that can be executed on a relational database (db). Specifically, given an nl and a db, the goal of Text-to-SQL is to generate an sql that accurately reflects the user’s intent and returns the appropriate results when executed on the database.
@@
Figure 1: Examples of the Text-to-SQL Task and Its Challenges.
II-B Text-to-SQL Human Workflow
When professional users (e.g., DBAs) perform the Text-to-SQL task, they first interpret the nl question, examine the database schema and contents, and then construct the corresponding sql based on their SQL expertise. Below, we outline this process in detail, as illustrated in Figure 1(a).
@@ II-B Text-to-SQL Human Workflow
Note that Step-2 aligns with the concepts of schema linking, database content retrieval, and additional information acquisition in recent Text-to-SQL solutions powered by language models (please refer to Figure 4 for more details).
Step-3: Writing SQL based on NL and DB Understanding: Finally, the DBA writes the corresponding sql based on the insights gained in Steps-1 and -2. This process, known as “Text-to-SQL Translation”, relies heavily on the DBA’s SQL expertise. However, this process can be very challenging due to the ambiguity of the nl or the complexity of the database. For example, as shown in Figure 1(a), despite understanding the need to link the Customer and Book tables, one must be familiar with the usage and norms of employing either a natural join or a subquery. In addition, there may be multiple possible sql queries because “genres” can refer to either “LiteraryGenre” or “SubjectGenre”.
Takeways. From the above steps, we intuitively identify three inherent challenges in the Text-to-SQL task: the uncertainty of the natural language, the complexity of the database, and the translation from the “free-form” natural language queries to the “constrained and formal” sql queries.
II-C Text-to-SQL Task Challenges
In this section, we will first discuss the fundamental challenges of the Text-to-SQL task. We will then analyze the technical challenges, i.e., the challenges we face when developing a strong Text-to-SQL solution in real-world scenarios.
C1: Uncertain Natural Language Query. Natural language queries often contain uncertainties due to ambiguity and underspecification [32]. In Text-to-SQL tasks, the challenges related to nl can be summarized as follows:
  • Lexical Ambiguity: This occurs when a single word has multiple meanings. For example, the word “bat” can refer to an animal or a baseball bat or the action of swinging.
@@ II-C Text-to-SQL Task Challenges
Figure 2: The Evolution of Text-to-SQL Solutions from the Perspective of Language Models.
C2: Complex Database and Dirty Content. The Text-to-SQL task requires a deep understanding of the database schema, including table names, columns, relationships, and data attributes. The complexity of modern schemas and large data volumes make this task especially challenging.
  • Complex Relationships Among Tables: Databases often contain hundreds of tables with complex interrelationships. Text-to-SQL solutions must accurately comprehend and leverage these relationships when generating sql.
  • Ambiguity in Attributes and Values: Ambiguous values and attributes in a database can complicate Text-to-SQL systems’ ability to identify the correct context.
  • Domain-Specific Schema Designs: Different domains often have unique database designs and schema patterns. The variations in schema design across domains make it difficult to develop a one-size-fits-all Text-to-SQL model.
@@
C3: Text-to-SQL Translation. The Text-to-SQL task differs from the compilation of a high-level programming language to a low-level machine language, as it usually has a one-to-many mapping between the input nl, db and output sql. Specifically, the Text-to-SQL task faces several unique challenges:
@@
  • Database Schema Dependency: The Text-to-SQL translation is highly dependent on the underlying database schema. As shown in Figure 1 (a) and (b), the same nl may produce different sql queries based on schema variations. This requires Text-to-SQL models to bridge gaps between training data and real-world schema differences.
Beyond the intrinsic challenges, developers must also overcome several technical obstacles to build reliable and efficient Text-to-SQL systems, as discussed below.
C4: Technical Challenges in Developing Text-to-SQL Solutions. Developing robust Text-to-SQL solutions requires addressing several key technical challenges, including:
  • Cost-effective Solution: Deploying Text-to-SQL models, particularly those using large language models, demands significant resources, such as hardware and/or API costs. Achieving an optimal balance between model performance and cost efficiency remains a crucial challenge.
@@
  • SQL Efficiency: The sql generated by Text-to-SQL models must be both correct and optimized for performance. This includes optimizing join operations, index usage, and query structures. Efficient queries reduce database load, improving system responsiveness and throughput.
  • Insufficient and Noisy Training Data: High-quality Text-to-SQL training data is challenging to obtain. Public datasets are often limited and may include noisy annotations, affecting model performance [33, 34]. Annotation requires database expertise, increasing costs, and the complexity of Text-to-SQL tasks often leads to errors.
  • Trustworthiness and Reliability: Text-to-SQL models must be trustworthy and reliable, consistently producing accurate results across diverse datasets and scenarios. Trustworthiness requires transparency, allowing users to understand and verify the generated sql.
@@ II-D Challenges Solving with Large Language Models
Difficulty Levels. We categorize the difficulty of Text-to-SQL into five levels, each addressing specific hurdles, as shown in Figure 2(a). The first three levels cover challenges that have been or are currently being addressed, highlighting the gradual progress in Text-to-SQL capabilities. The fourth level includes challenges that are the focus of current LLM-based solutions, while the fifth level represents future challenges, showing our vision for Text-to-SQL advancements over the next five years.
The Evolution of Text-to-SQL Solutions. The development of Text-to-SQL solutions, illustrated in Figure 2(b), progresses through four distinct stages: the rule-based stage, the neural network-based stage, the PLM-based stage, and the LLM-based stage. At each stage, we analyze shifts in target users, i.e., from experts to broader user groups, and the extent to which various Text-to-SQL challenges are addressed.
Figure 3: The Categorization of PLM and LLM in Text-to-SQL.